OpenStudy (anonymous):

I have worked on this problem for 2 hours. I have exhausted all but one last chance to answer it. The problem is as follows: Expensive amplifier A is rated at 260W , while the more modest amplifier B is rated at 46W . Estimate the sound level in decibels you would expect at a point 3.5m from a loudspeaker connected in turn to each amp. My answers are not right, but the program is telling me I'm close. I submitted 118 and 134. PLEASE HELP! <3 Thank you in advance. And please show me how.

OpenStudy (anonymous):

\[P_{dB}=120+10\log_{10}\left( P_{\frac{ W }{ m^2 }} \right)=120+10\log_{10}\left( \frac{ P_W}{4 \pi d^2 } \right)\]

OpenStudy (anonymous):

For amplifier 1, I get 122.28dB and for amplifier 2, 114.75dB The difference is 122.28-114.75=7.52dB=10log(P1/P2)=10log(260/46)

OpenStudy (anonymous):

THANK YOU!!!!!